# The Learning Approach to Games

### v1

@@ Abstract
This work provides a unified framework for exploring games. In existing literature, strategies of players are typically assigned scalar values, and the concept of Nash equilibrium is used to identify compatible strategies. However, this approach lacks the internal structure of a player, thereby failing to accurately model observed behaviors in reality. To address this limitation, we propose to characterize players by their learning algorithms, and as their estimations intrinsically induce a distribution over strategies, we introduced the notion of equilibrium in terms of characterizing the recurrent behaviors of the learning algorithms. This approach allows for a more nuanced understanding of players, and brings the focus to the challenge of learning that players face. While our explorations in discrete games, mean-field games, and reinforcement learning demonstrate the framework’s broad applicability, they also set the stage for future research aimed at specific applications.
@@ 1 Introduction
Note that a common domain ${\cal D}$ is not restrictive and might have subdomains that each estimation depends on. As an example, it is standard to have a model with time and state space included in ${\cal D}$. Also, tpically ${\cal E}\subset\mathbb{A}\times{\cal D}$. When there are more than one player, we keep track of them by the index $i\in\mathbb{N}_{0}:=\{1,2,\dots\}$, and set $\vec{\varphi}:=\prod_{i\in\mathbb{N}_{0}}\varphi^{i}$ for any notation in the Definition 1. In this case, typically ${\cal E}^{i}\subset\vec{\mathbb{A}}\times\vec{\cal D}$ and players are observing each other. As an example, there might be a common observed state space, or players might observe each other’s actions.
@@
We can generalize the definition to more than one player in a straightforward manner. In this case, as they interact with each other, a collectively recurrent behavior forms the basis for an equilibrium. However, to improve the characterization of an equilibrium, one needs to further empose conditions on the estimations too. Throughout this work, we will explore and modify these definitions to settings that are commonly considered in the literature.
@@ 2 Discrete Games
A crucial notion to introduce is the value of a player. We observe that value is fundamentally a formalism used to determine an optimal action. We do not regard it as something tangible that a player necessarily obtains, which might only be the case in limited situations. However, even in these situations, value is typically considered an expectation rather than a definitive quantity. Now, we introduce two concepts: transition costs and state values, which are akin to running costs and terminal costs in the literature [^3] We use cost and value interchangebly. In the case of scalar objectives as in this work, distinction is more pronounced. However, for multi-objective frameworks, there is typically no binary choice, but rather a continuum of choices.:
@@
which is a random variable on $\hat{\Omega}$. Set $\mathscr{M}_{J}^{i}$ as the space of all such functions $(\hat{\Omega}\times\mathbb{T}\times\Omega\times{\cal A}^{i}\to\mathbb{R})$.
@@
which is more intuitive and useful in computations, but we won’t keep track of it. Moreover, one can further extend the space of parameters. Simplest example might be to write $\hat{F}(t,s,x,a)$. We are trying to point out various modeling choices that we made to cover enough cases to promote our framework; however, it is necessary to adjust these appropriately based on the problem under consideration.
@@
Note that, in general, the equality should be understood in terms of two subsets of controls having equal probabilities, when they are equal once extended with the relation $=^{t,x;i}$. Also, we didn’t suppress $\hat{\omega}$ as it will be integrated. Here, at $(\hat{\omega},t,x)$, $\hat{\pi}$ approximates the potential optimal controls for $J(\hat{\omega},t,x,\cdot)$, which will be dictated by the equilibrium condition below.
@@
Although apriori ${\Upsilon}^{t,x;i}$ appears to be solely induced by $\hat{\pi}$, $\hat{\pi}$ itself is a function of the value $J^{i}$, and hence ${\Upsilon}^{t,x;i}$ is a function from $\mathscr{M}_{T}\times\mathscr{M}_{p}\times\mathscr{M}_{\Gamma}\times\mathscr{M}_{F}\times\mathscr{M}_{\phi}\times\mathscr{M}_{\pi}$. In the two-player example discussed in section 2.1, $\hat{\pi}$ solves the optimization by brute force, yielding a deterministic action for each scenario in $\hat{\Omega}$. In the control problem described in section 4, $\hat{\pi}$ constructs the distribution over strategies relying again on the associated values. Let us note that we induce a distribution over controls because it is a more familiar and convenient choice; however, by considering sequences of future states, one can easily induce a distribution over sequences of actions. Moreover, this behavior, represented by ${\Upsilon}$, is well-suited for the learning or playing phase. One might want to modify the behavior of the player to yield the most likely control during the competition phase.
It is important to motivate the role of randomness in value, which then induces a probability distribution over actions in (2.8). Recall that in sufficiently complex settings, such as chess, values are inherently unknown and must be learned through significant effort. That is, the randomness of the value models what is unknown to the player. The key role of randomness in value is to allow players to explore systematically. If the player is not satisfied with the current estimates, it is natural to assign greater probabilities to unexplored controls for having higher values. This approach naturally leads to the search for controls with more satisfactory results. We will demonstrate a toy version for the two-player game in Section 2.1. Consider a daily life scenario: suppose you have a favorite dish, and your friend suggests a new one you have never tried. Since you expect it is not better, if your model only considers expectations, you won’t explore. However, if you model the value as random, you might try the new dish. This approach aligns with the common intuition that a better understanding of values should lead to less uncertain strategies. Once you try the new dish, you know its value to you more accurately, and that experience might deter further exploration. Again, the design of a player is a crucial aspect of understanding games.
@@
and call ${}^{0}\hat{\varphi}^{i}$ as the prior of player $i$. Note that, $({}^{n}\hat{T}^{i},{}^{n}\hat{p}^{i},{}^{n}\hat{\Gamma}^{i},{}^{n}\hat{F}^{i},{}^{n}\hat{\phi}^{i})$ defines ${}^{n}J_{\mathfrak{L}}^{i}(t,x;\alpha)$ as in (2.6), which then together with ${}^{n}\hat{\pi}^{i}$ induces ${\Upsilon}_{\mathfrak{L}}^{t,x;i}$ on ${}^{n}{\cal A}^{t,x;i}$ as in (2.8) and ${}^{n}\gamma_{\mathfrak{L}}^{t,x;i}$ on $\mathbb{A}^{t,x;i}$ as in (2.9). [^4] We may drop $\mathfrak{L}$ from subscripts if it is clear from the context.
We say all players are behaving rationally, if at each $(t,x)\in\mathbb{T}\times\mathbb{S}$, they draw an action from the distribution ${}^{n}\vec{\gamma}_{\mathfrak{L}}^{t,x}:=({}^{n}\gamma_{\mathfrak{L}}^{t,x;1},{}^{n}\gamma_{\mathfrak{L}}^{t,x;2},\dots)$, or draw from the distribution ${}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x}:=({}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;1},{}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;2},\dots)$ to decide on strategies for $\mathbb{T}_{t+\hat{T}}^{t}$. We always assume that players are learning under rational behaviors. In a two-player game, for example, if one player behaves nonsensically, expecting any type of convergent behavior becomes meaningless. We emphasize that this assumption of rationality is not a constraint but rather means that we model players within the current framework of this work.
@@
As a first example, suppose there exists finitely many models $\{\vec{\varphi}_{m}\}_{m=1}^{M}$ for which $\vec{\mathfrak{L}}_{\varphi}$ can take values in. Furthermore,
$\mathbb{P}^{u}\big{(}{}^{n+1}\vec{\mathfrak{L}}_{\varphi}=\vec{\varphi}_{m}|\ {}^{n}\vec{\mathfrak{L}}_{\varphi}=\vec{\varphi}_{\ell}\big{)}>0,\quad\forall\ell,m\in\{1,\dots,M\},\ n\in\mathbb{N},\varphi\in\{T,p,\Gamma,F,\phi,\pi\}$
then, all of the models satisfy (iii) in the definition of uncertain equilibrium. As long as players are able to solve their own optimization problem, all the models are uncertain equilibrium. Of course, in reality it is unlikely to have finitely many potential models. Basic games like rock-paper-scissors can be modeled in this way. Here, only relevant model is about what the other player will do, that is $\Gamma^{1},\Gamma^{2}$. Then by construction, we can simplify to only $3$ different models assigning certain actions is enough. Then, depending on the ongoing observations, players will continuously update their estimations on $\Gamma$ and act accordingly. [^5] This may not be a natural way to model, but rather to provide a straightforward example.
@@
To briefly elaborate on how players solve their own optimization problems, this is partly achieved by repeating past observations with evolving estimations. As a player learns, i.e. as $n$ increases, and recalls past observations, the exploration of other potential scenarios under ${}^{n}\hat{\varphi}^{i}$ aims to capture the term $\sup_{\alpha\in\mathcal{A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\alpha)$. During this revaluation and exploration process, new strategies may be discovered, or a new assessment of value might lead to changes in ${}^{n}\hat{\pi}^{i}$. Condition (ii) in the definition of uncertain equilibrium implies that the player has explored potential strategies and is capable of generating the best ones under various scenarios of $\hat{\Omega}$ as learning continues.
On the other hand, the values of the player are driven externally. There is no universal concept of what holds higher value; rather, these concepts are shaped by factors such as needs, interactions, and self-evaluations, as evidenced by the highly diverse values found across individuals and societies. We would like to emphasize that the equilibrium can be understood specifically for each $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$. With this context in mind, as the player considers a past state that could potentially recur, condition (iii) in the definition of uncertain equilibrium implies that the induced distribution over strategies also recurs as learning progresses.
The choice of $\operatorname*{liminf}$ instead of $\operatorname*{limsup}$ is important. Let us consider two scenarios:
  • (i)
    Players might remain in a particular equilibrium for nearly all of the time as learning progresses, but they may occasionally explore other actions. Such behavior is observable even in the simplest settings, as we will demonstrate later. We still recognize this as an equilibrium. The use of $\operatorname*{liminf}$ implies that the players are within an $r$-distance of this particular equilibrium infinitely often.
  • (ii)
    There might be two or more equilibria between which players switch. We recognize each one as an equilibrium, as players will find themselves in each particular equilibrium infinitely often. If we were to require $\operatorname*{limsup}$ instead, it would imply that players are stuck in a particular equilibrium indefinitely, which is quite hard to expect in reality.
@@
Recall that if regret is defined similarly to the correlated equilibrium and is assumed to be sublinear, then by definition, the observed time-averaged joint distribution over actions converges to a correlated equilibrium. Considering the estimated distributions over strategies in (2.12), one might expect that all players will eventually induce the same distribution in symmetric situations. However, in general, there is no reason to assume that a single distribution will characterize every player’s considerations.
The concept of Nash equilibrium focuses solely on controls, according to their associated scalar values, inherently excluding the intrinsic structure of a player. In situations where a central planner announces policies for individual agents, such as environmental regulations, traffic management, public health initiatives, with the knowledge that every individual will act according to their own assesment of value, Nash equilibrium is definitely the right consideration. The central planner needs to model agent’s individual values to construct stable policies. In such cases, it is not meaningful to model each and every agent in detail of their learning algorithms. We also remark that, the Nash equilibrium requires players to have exact knowledge of the strategies of other players. In a pure equilibrium sense, this instability is significant enough that an equilibrium does not exist even for simplest games like rock-paper-scissors. To address this, one typically adopts relaxed controls, which is essential but still omits details about the underlying player.
@@
To relate this quantity to familiar concepts with which we all can relate, we say that at the state $(t,x)\in\mathbb{T}\times\Omega$, the player $i$ is currently desperate if ${}^{n}\kappa^{i}(t,x)=0$, and euphoric if ${}^{n}\kappa^{i}(t,x)=1$. If the player is desperate, as the learning progresses, either ${}^{n}\hat{B}^{i}$ will decrease, leading the player, in some sense, to accept the situation; or ${}^{n}J^{i}$ will assign higher values to underexplored strategies, encouraging the player to explore them. Let us note that it is not necessarily better for the learning algorithm to adjust ${}^{n}J^{i}$, since ${}^{n}\hat{B}^{i}$ might be unrealistically high. One can describe the player’s current situation in more detail by using verbal subcategories like
@@
Next, time-consistency, or Dynamic Programming Principle, can be naturally introduced for any learning parameter. The most important and familiar one is for the value function, and it also relates $\hat{\phi}$ to the concept of the standard value function in the literature. We say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time consistent value almost surely, if for any $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$ and $t\leq T_{0}\leq t+\hat{T}(t,x)$, it holds
@@
which closely resembles the standard time-consistency, or Dynamic Programming Principle, for the standard value function. Similarly, for example, we say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time-consistent distribution over controls almost surely at $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$, if for any $t\leq T_{0}\leq t+\hat{T}(t,x)-1$, and $x_{T_{0}}\in\mathbb{S}_{T_{0}}$, $\hat{\pi}(\hat{\omega},t,x)$ induces the same distribution as $\hat{\pi}(\hat{\omega},T_{0},x_{T_{0}})$ on the space ${\cal A}^{\hat{T}(t,x);T_{0},x_{T_{0}},i}$. Here, the quotient space is defined similarly, with the relation terminating at $t+\hat{T}(t,x)-1$ instead of $t+\hat{T}(T_{0},x_{T_{0}})-1$.

### v2

@@ Abstract
This work introduces a unified framework for a more detailed exploration of games. In existing literature, strategies of players are typically assigned scalar values, and the concept of Nash equilibrium is used to identify compatible strategies. However, this approach lacks the internal structure of a player, thereby failing to accurately model observed behaviors.
To address this limitation, we propose an abstract definition of a player. This allows for a more nuanced understanding of players and brings the focus to the challenge of learning that players face. Unlike Markov decision processes, which formalize control problems but not agent design, our framework subsumes standard reinforcement learning structures. It thus offers a language that enables a deeper connection between games and learning. To illustrate the need for such generality, we study a simple two-player game and show that even in the most basic settings, a sophisticated player may adopt dynamic strategies that cannot be captured by simpler designs or compatibility analysis alone.
In the discrete setting, we consider a player whose structure incorporates standard estimates from the literature. We explore connections to correlated equilibrium and highlight that dynamic programming naturally applies to all estimates. In the mean-field setting, we exploit symmetry to construct explicit examples of equilibria. Finally, we examine connections to reinforcement learning and bandit problems, demonstrating the broad applicability of the framework.
ACM Classification: I.2.6; J.4
@@ 1 Introduction
To motivate our goal intuitively, consider a swarm of self-driving cars operating in a city. From the traditional game-theoretic viewpoint, one might analyze the density of cars on the streets and emerging traffic patterns. For such a problem, it is neither feasible nor meaningful to account for the detailed internal structure of each vehicle. As a result, however, it cannot be used to specify or simulate how the cars themselves make decisions. We remark that such traditional methods remain rooted in a central-planner perspective. Roughly speaking, equilibrium concepts involve solving a control problem, augmented by a compatibility condition among agents’ strategies. When the goal is to design optimal policies within well-defined environments, the traditional framework remains appropriate. However, when the focus shifts to designing complex players themselves, such as those enabled by modern reinforcement learning, the available methods are powerful, but a unifying language is lacking. This work aims to fill this gap.
@@
Note that a common domain ${\cal D}$ is not restrictive and might have subdomains that each estimation depends on. As an example, it is standard to have a model with time and state space included in ${\cal D}$. Also, typically ${\cal E}\subset\mathbb{A}\times{\cal D}$. When there are more than one player, we keep track of them by the index $i\in\mathbb{N}_{0}:=\{1,2,\dots\}$, and set $\vec{\varphi}:=\prod_{i\in\mathbb{N}_{0}}\varphi^{i}$ for any notation in the Definition 1. In this case, typically ${\cal E}^{i}\subset\vec{\mathbb{A}}\times\vec{\cal D}$ and players are observing each other. As an example, there might be a common observed state space, or players might observe each other’s actions.
@@
We can generalize the definition to more than one player in a straightforward manner. In this case, as they interact with each other, a collectively recurrent behavior forms the basis for an equilibrium.
@@
The main objective of this work is to show that Definition 1 is sufficiently general to encompass any player design found in the literature. A more important claim is that observed player behavior must be understood via conditions on $({\cal O},{\cal L}_{1},\dots,{\cal L}_{k},{\Upsilon})$. It is clear that some conditions on players’ observations must hold, and this is dictated by the model at hand. On the other hand, Definition 2 provides a natural condition requiring the recurrence of observed behaviors. After introducing specific estimations, we discuss how to impose conditions such as optimality, consistency, and even psychological states to better characterize both the players and the equilibrium concept.
@@
Literature Review. Classic non-cooperative game theory traces back to the von Neumann-Morgenstern seminal work [28] and Nash [27]. We refer to the book Maschler-Solan-Zamir [24] for a comprehensive treatment. There are many notions of equilibrium, and we highlight the correlated equilibrium introduced by Aumann [2]. The author raises common criticisms of the classic Nash equilibrium and introduces correlated equilibrium, which incorporates randomness in players’ strategies. A correlated equilibrium is a single distribution over the players’ strategies and is therefore typically motivated by a mediator who draws from this distribution. However, a single distribution cannot capture the differing knowledge each player may possess. Because of this structural connection, we will offer a comparison after introducing the uncertain equilibrium for discrete games.
Learning in games has a rich literature, beginning with Brown [7], who introduced the notion of fictitious play. Players are assumed to have predefined learning rules, and the question is whether the long-run average of observed actions converges to an equilibrium. Although such convergence is not always guaranteed (see, for example, Daskalakis et al. [15]), Hart and Mas-Colell [17] combine regret with fictitious play to show convergence to correlated equilibrium. For a comprehensive treatment, see the book Fudenberg-Levine [16].
To address games with a large number of players, where equilibria become intractable, Lasry-Lions [22] and, independently, Huang-Malhamé-Caines [18] introduced the concept of mean-field games. Since then, the framework has been extensively studied. In this setting, agents interact only through the empirical distribution of their states and are indistinguishable from one another, allowing a continuum limit to be identified with a representative agent. We refer the reader to the excellent two-volume book Carmona-Delarue [12, 13].
In our framework, players might favor having a large number of estimates, and each of them might need to be learned from observations. We cannot hope to cover every relevant learning algorithm and its literature; however, we point out the diverse work on random value functions. We also advocate that the value is fundamentally unknown to a learning player and is one of the main sources of uncertainty in planning future behavior. We refer to Thurstone [31], Luce [23], Block [6], McFadden [25], Train [32], and references therein for discussions rooted in psychology and economics. One of the oldest approaches is Thompson sampling, introduced in Thompson [30] and recently popularized by the empirical study Chapelle-Li [14]. Upper confidence bound algorithms in the context of multi-armed bandit problems, see for example Auer et al. [1], can also be viewed as a random value approach. Lastly, let us mention Bellemare-Dabney-Munos [5] who prominently promoted modeling the value function as a random variable in the context of Markov decision processes with applications to reinforcement learning.
@@ 2 Discrete Games
A crucial notion to introduce is the value of a player. We observe that value is fundamentally a formalism used to determine an optimal action. We do not regard it as something tangible that a player necessarily obtains, which might only be the case in limited situations. However, even in these situations, value is typically considered an expectation rather than a definitive quantity. Now, we introduce two concepts: transition costs and state values, which are akin to running costs and terminal costs in the literature [^3] We use cost and value interchangeably. In the case of scalar objectives as in this work, distinction is more pronounced. However, for multi-objective frameworks, there is typically no binary choice, but rather a continuum of choices.:
@@
which is a random variable on $\hat{\Omega}$. Set $\mathscr{M}_{J}^{i}$ as the space of all such functions $(\hat{\Omega}\times\mathbb{T}\times\Omega\times{\cal A}^{i}\to\mathbb{R})$. We point out that there is no fundamental difference between requiring either a distribution or a random variable, but the latter allows us to refer to samples.
@@
which is more intuitive and useful for computations, but we will not keep track of it. Moreover, one can further extend the space of parameters. A simple example would be to write $\hat{F}(t,s,x,a)$. We are merely pointing out basic modeling choices, with the understanding that these will need to be tailored to the specific problem at hand. In fact, a truly complex player, such as a human, would not only adapt existing estimates and behaviors, but also develop new types of estimates and behaviors over time. Just as a child starts with limited capacities and gradually acquires a rich repertoire, such dynamic structural growth might indeed be essential for modeling higher-order intelligence. While this remains far beyond current work, it represents a compelling direction for future research.
@@
Here, at $(\hat{\omega},t,x)$, $\hat{\pi}$ approximates the potential optimal controls for $J(\hat{\omega},t,x,\cdot)$, which will be dictated by the equilibrium condition below. Note that, the equality should be understood in terms of two subsets of controls having equal probabilities, when they are equal once extended with the relation $=^{t,x;i}$. Also, we didn’t suppress $\hat{\omega}$ as it will be integrated.
@@
Although apriori ${\Upsilon}^{t,x;i}$ appears to be solely induced by $\hat{\pi}$, $\hat{\pi}$ itself is a function of the value $J^{i}$, and hence ${\Upsilon}^{t,x;i}$ is a function from $\mathscr{M}_{T}\times\mathscr{M}_{p}\times\mathscr{M}_{\Gamma}\times\mathscr{M}_{F}\times\mathscr{M}_{\phi}\times\mathscr{M}_{\pi}$. We introduced $\hat{\pi}$ to carry out the discussions more explicitly. [^4] When $\hat{\pi}$ is a determined function of the underlying estimates, as in the case of our examples in this work, it is more appropriate to embed it into the behavior ${\Upsilon}$. In the case where strategies are learned both from observations and other estimates, it is even more favorable to adjust the Definition 1 slightly so that the domains can include spaces of estimations. We avoid this, as it would not enhance the discussion but would complicate the notation. In the two-player example discussed in section 2.1, $\hat{\pi}$ solves the optimization by brute force, yielding a deterministic action for each scenario in $\hat{\Omega}$. In the control problem described in section 4, $\hat{\pi}$ constructs the distribution over strategies relying again on the associated values. Let us note that we induce a distribution over controls because it is a more familiar and convenient choice; however, by considering sequences of future states, one can easily induce a distribution over sequences of actions.
This behavior, represented by $\hat{\pi}$, is well suited for the learning or playing phase and is, in essence, Thompson sampling adapted to our framework. During a competition phase, one might modify $\hat{\pi}$ to be deterministic, selecting the control corresponding to the mode of the value. In particular situations, such as performing surgery, it is not only wrong but also unethical to forgo the most likely action and instead select one at random. Likewise, a deterministic $\hat{\pi}$ might instead choose the control that maximises a confidence bound of the value. Moreover, a player may adjust its behavior to occasionally select unlikely actions, exploring states that are disadvantageous or even entirely unseen when facing a weak opponent.
It is important to motivate the role of randomness in value, which then induces a probability distribution over actions in (2.8). Recall that in sufficiently complex settings, such as chess, values are inherently unknown and must be learned through significant effort. That is, the randomness of the value models what is unknown to the player. The key role of randomness in value is to allow players to explore systematically. If the player is not satisfied with the current estimates, it is natural to assign greater probabilities to unexplored controls for having higher values. This approach naturally leads to the search for controls with more satisfactory results. We will demonstrate a toy version for the two-player game in Section 2.1. Consider a daily life scenario: suppose you have a favorite dish, and your friend suggests a new one you have never tried. Since you expect it is not better, if your model only considers expectations, you won’t explore. However, if you model the value as random, you might try the new dish. This approach aligns with the common intuition that a better understanding of values should lead to less uncertain strategies. Once you try the new dish, you know its value to you more accurately, and that experience might deter further exploration.
@@
and call ${}^{0}\hat{\varphi}^{i}$ as the prior of player $i$. Note that, $({}^{n}\hat{T}^{i},{}^{n}\hat{p}^{i},{}^{n}\hat{\Gamma}^{i},{}^{n}\hat{F}^{i},{}^{n}\hat{\phi}^{i})$ defines ${}^{n}J_{\mathfrak{L}}^{i}(t,x;\alpha)$ as in (2.6), which then together with ${}^{n}\hat{\pi}^{i}$ induces ${\Upsilon}_{\mathfrak{L}}^{t,x;i}$ on ${}^{n}{\cal A}^{t,x;i}$ as in (2.8) and ${}^{n}\gamma_{\mathfrak{L}}^{t,x;i}$ on $\mathbb{A}^{t,x;i}$ as in (2.9). [^5] We will drop $\mathfrak{L}$ from subscripts if it is clear from the context.
We say all players are behaving rationally, if at each $(t,x)\in\mathbb{T}\times\mathbb{S}$, they draw an action from the distribution ${}^{n}\vec{\gamma}_{\mathfrak{L}}^{t,x}:=({}^{n}\gamma_{\mathfrak{L}}^{t,x;1},{}^{n}\gamma_{\mathfrak{L}}^{t,x;2},\dots)$, or draw from the distribution ${}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x}:=({}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;1},{}^{n}{\Upsilon}_{\mathfrak{L}}^{t,x;2},\dots)$ to decide on strategies for $\mathbb{T}_{t+\hat{T}}^{t}$. [^6] These are equivalent under time-consistency. The concept is discussed later in this section. We always assume that players are learning under rational behaviors. In a two-player game, for example, if one player behaves nonsensically, expecting any type of convergent behavior becomes meaningless. We emphasize that this assumption of rationality is not a constraint but rather means that we model players within the current framework of this work.
@@
To briefly elaborate on how players solve their own optimization problems, this is partly achieved by repeating past observations with evolving estimations. As player learns, that is as $n$ increases, and recalls past observations, the exploration of other potential scenarios under the current estimations ${}^{n}\hat{\varphi}^{i}$ aims to capture the term $\sup_{\alpha\in\mathcal{A}^{i}}{}^{n}J^{i}(\hat{\omega},t,x,\alpha)$. During this revaluation and exploration process, new strategies may be discovered, or a new assessment of value might lead to changes in ${}^{n}\hat{\pi}^{i}$. Condition (ii) in the definition of uncertain equilibrium implies that the player has explored potential strategies and is capable of generating the best ones under various scenarios of $\hat{\Omega}$ as learning continues.
On the other hand, the values of the player are driven externally. Not only is there no universal notion of what holds higher value, but values often involve multiple, conflicting objectives. These are shaped by needs, interactions, and self-evaluations, as reflected in the remarkable diversity of values across individuals and societies. We would like to emphasize that the equilibrium can be understood specifically for each $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$. With this context in mind, as the player considers a past state that could potentially recur, condition (iii) in the definition of uncertain equilibrium implies that the induced distribution over strategies also recurs as learning progresses.
The choice of $\operatorname*{liminf}$ instead of $\operatorname*{limsup}$ is important. Let us consider two scenarios: (i) Players might remain in a particular equilibrium for nearly all of the time as learning progresses, but they may occasionally explore other actions. Such behavior is observable even in the simplest settings, as we will demonstrate later. We still recognize this as an equilibrium. The use of $\operatorname*{liminf}$ implies that the players are within an $r$-distance of this particular equilibrium infinitely often. (ii) There might be two or more equilibria between which players switch. We recognize each one as an equilibrium, as players will find themselves in each particular equilibrium infinitely often. If we were to require $\operatorname*{limsup}$ instead, it would imply that players are stuck in a particular equilibrium indefinitely, which is quite hard to expect in reality.
@@
Recall that if regret is defined similarly to the correlated equilibrium and is assumed to be sublinear, then the observed time-averaged joint distribution over actions converges to a correlated equilibrium. Considering the estimated distributions over strategies in (2.12), one might expect that all players will eventually induce the same distribution in symmetric situations. However, in general, there is no reason to assume that a single distribution will characterize every player’s considerations.
The concept of Nash equilibrium focuses solely on controls, according to their associated scalar values, inherently excluding the intrinsic structure of a player. For example, in situations where a central planner announces policies for individual agents, such as environmental regulations, traffic management, public health initiatives, with the knowledge that every individual will act according to their own assessment of value, Nash equilibrium is definitely the right consideration. The central planner needs to model agent’s individual values to construct stable policies. In such cases, it is not meaningful to model each and every agent in detail of their learning algorithms. We also remark that, the Nash equilibrium requires players to have exact knowledge of the strategies of other players. In a pure equilibrium sense, this instability is significant enough that an equilibrium does not exist even for simplest games like rock-paper-scissors. To address this, one typically adopts relaxed controls, which is essential but still omits details about the underlying player.
@@
To relate this quantity to familiar concepts with which we all can relate, we say that at the state $(t,x)\in\mathbb{T}\times\Omega$, the player $i$ is currently desperate if ${}^{n}\kappa^{i}(t,x)=0$, and euphoric if ${}^{n}\kappa^{i}(t,x)=1$. If the player is desperate, as the learning progresses, either ${}^{n}\hat{B}^{i}$ will decrease, leading the player, in some sense, to accept the situation; or ${}^{n}J^{i}$ will assign higher values to underexplored strategies, encouraging the player to explore them. One can describe the player’s current situation in more detail by using verbal subcategories like
@@
Next, time-consistency, or Dynamic Programming Principle, can be naturally introduced for any learning parameter and can likewise be required at equilibrium. The most important and familiar one is for the value function, and it also relates $\hat{\phi}$ to the concept of the standard value function in the literature. We say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time consistent value almost surely, if for any $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$ and $0\leq T_{0}\leq\hat{T}(t,x)$, it holds
@@
which closely resembles the standard time-consistency, or Dynamic Programming Principle, for the standard value function.
To illustrate how time-consistency can be required of an estimation, we focus on $\hat{\pi}$; similar conditions can be formulated for $\hat{T}$, $\hat{p}$, and $\hat{\Gamma}$ as well. We say that the estimate $(\hat{T},\hat{p},\hat{\Gamma},\hat{F},\hat{\phi},\hat{\pi})$ yields a time-consistent distribution over controls almost surely at $(t,x)\in\mathbb{T}\times\mathbb{S}_{t}$, if for any $t\leq T_{0}\leq t+\hat{T}(t,x)-1$, and $x_{T_{0}}\in\mathbb{S}_{T_{0}}$, $\hat{\pi}(\hat{\omega},t,x)$ induces the same distribution as $\hat{\pi}(\hat{\omega},T_{0},x_{T_{0}})$ on the space ${\cal A}^{\hat{T}(t,x);T_{0},x_{T_{0}},i}$. Here, the quotient space is defined similarly, with the relation terminating at $t+\hat{T}(t,x)-1$ instead of $t+\hat{T}(T_{0},x_{T_{0}})-1$.